OpenStudy (anonymous):

Differentiate the function. F(y) = y ln(6 + ey)

OpenStudy (anonymous):

\[F(y)=y\ln(6+e^y)~?\]

OpenStudy (anonymous):

they said that is incorrect

zepdrix (zepdrix):

\[\large F'(y)=\color{royalblue}{(y)'}\ln(6+e^y)+y\color{royalblue}{\left[\ln(6+e^y)\right]'}\] So it looks like you'll need to apply the product rule.

OpenStudy (anonymous):

@jkool99, I was merely asking for confirmation. Is that the same as your given function? Your writing is a bit ambiguous.

OpenStudy (anonymous):

lol i completely missed read your response. yeah that is the problem what you wrote. i thought that was the solution that was given

OpenStudy (anonymous):

i got it now though. thanks @zepdrix